OpenStudy (adi3):

please help will medal For each of the following functions: (i) find f^-1 (x) and the funtions are f(x) = (3 - 2x)/4

OpenStudy (adi3):

We need to inverse \[\frac{ 3-2x }{ 4}\]

OpenStudy (adi3):

but I don't know how to fo it.

OpenStudy (adi3):

@phi

OpenStudy (adi3):

I don't get that

rishavraj (rishavraj):

replace f(x) with "y" so its y = (3 - 2x)/4

rishavraj (rishavraj):

and then switch x and y .... so u will get x = (3 - 2y)/4 and then solve for y

OpenStudy (adi3):

so for solving for y we do